OpenStudy (anonymous):

i solved for 3x+2y=0 but not sure wich one of these is correct? (0,0), (0-3), (-2,3) or (0,0), (0,-3), (-2,0)

OpenStudy (aravindg):

just plug in the value and see

OpenStudy (aravindg):

what do you get when you plug in (-2,3) ?

OpenStudy (anonymous):

-6

Directrix (directrix):

plug in (-2,3) ------------- 3x+2y=0 3(-2) + 2(3) = ? Hint: It is not -6.

OpenStudy (anonymous):

0

Directrix (directrix):

Correct. So, which of these is the answer you seek: i solved for 3x+2y=0 but not sure wich one of these is correct? (0,0), (0-3), (-2,3) @krissywatts

OpenStudy (anonymous):

(0,0) ( 2,-3) (-2,3) is this correct?

Directrix (directrix):

@krissywatts Exactly ONE of those three ordered pairs is correct. You have already figured it out. Click on the attachment to see your work.

OpenStudy (anonymous):

i have to find three

Directrix (directrix):

(0,0) ( 2,-3) (-2,3) is this correct? YES, these three are correct. @krissywatts

OpenStudy (anonymous):

okay thanks